Sue has $3.50 of dimes and nickels. If she has 50 coins, how many of each type of coin does she have?
Answer by checkley71(8403)

350=10X+5(50-X) OR 350=10X+250-5X OR 100=5X OR X-100/5 OR X=20 DIMES THUS WE HAVE 50-20=30 NICKELS.
PROOF 10*20+5*30=350 OR 200+150=350 OR 350=350